OpenStudy (anonymous):

12y-8 over -4? anyone?? We have to simplify the expression.

OpenStudy (anonymous):

y=-2/12

OpenStudy (danjs):

(12y-8) / - 4 or 12y - (8/-4)

OpenStudy (danjs):

group multiple term numerator and denominator with parenthesis

OpenStudy (anonymous):

I meant 3y+2

OpenStudy (danjs):

remember the properties of addition of fractions \[\frac{ 12y-8 }{ -4 } = \frac{ 12y }{ -4 }-\frac{ 8 }{ -4 }\]

OpenStudy (danjs):

simplify both those fractions if possible

OpenStudy (anonymous):

ok

OpenStudy (danjs):

(12/-4) *y - (8/(-4)) = -3y - (-2) = -3y+2 close

OpenStudy (drigobri):

\[\frac{ 12y - 8 }{ -4 } = \frac{ 12y}{ -4 } + \frac{- 8 }{ -4 } = -3y + 2\] If you want find the Y: \[-3y + 2 = 0\]\[3y = 2\]\[y = \frac{ 2 }{ 3 }\]\[y = 1,5\]

OpenStudy (danjs):

You can not solve for anything, How do you figure you just set that = to zero?

OpenStudy (danjs):

-3y + 2 nothing to solve , simplified

OpenStudy (drigobri):

Hey my friend, i just said that if he want to find the Y, he can set -3y + 2 = 0 why you so angry? lol